# The Three Hundred Project: deducing the stellar splashback structure of galaxy clusters from their orbiting profiles

### v1

@@ 6 Conclusions
    Sharper Stellar Truncation: The stellar orbiting profile exhibits a steeper truncation than the dark matter, consistent with a sharper splashback feature. This is in line with the findings of Deason et al. (2021) and O’Neil et al. (2021), and suggests that stellar material - in particular the ICL - may offer a more distinct observational signature of the splashback boundary.
@@
    Accretion Rate Dependence: Both components follow the same $r_{\rm t}-\Gamma$ relation, enabling the use of stellar profiles as tracers of halo mass accretion history. This builds on the theoretical framework established by Adhikari et al. (2014), More et al. (2015), and Diemer (2017), and confirms that the splashback radius is sensitive to recent mass growth, with minimal dependence on halo mass or cosmology.
@@
    Observational Viability: Projected stellar density profiles allow recovery of the splashback radius $R_{\rm t}$ with modest scatter ($\sim 0.3R_{200m}$), demonstrating its observational viability. This complements observational detections of splashback features in galaxy density and weak lensing profiles (e.g. Chang et al. 2018b; Baxter et al. 2017; Gabriel-Silva & Sodré 2025) and supports the use of the ICL as a tracer of cluster boundaries, especially in deep imaging surveys.
@@
    Infalling and Inner Profile Correlations: The infalling and inner stellar profiles correlate with those of the dark matter, particularly in non-parametric measures such as the scale radius $r_{s}$. This suggests that stellar material can be used to probe both the outer and inner dynamical structure of clusters, consistent with results from Pizzardo et al. (2024) and O’Neil et al. (2021).
@@
As upcoming surveys like Euclid, LSST, and Roman deliver deeper imaging and improved statistics, the stellar splashback radius may become a key tool in precision cluster cosmology. For this reason, there should be a focus on developing observational pipelines for ICL-based splashback detection, quantifying systematics due to projection and substructure, and integrating splashback measurements into multi-probe cosmological frameworks (e.g. Mpetha et al., 2025; Joshi et al., 2025).
@@ Acknowledgements
KW and CP acknowledge the support of the ARC Centre of Excellence for All Sky Astrophysics in 3 Dimensions (ASTRO 3D), through project number CE170100013. AK is supported by the Ministerio de Ciencia e Innovación (MICINN) under research grant PID2021-122603NB-C21 and further thanks Hüsker Dü for new day rising. WC is supported by the STFC AGP Grant ST/V000594/1, the Atracción de Talento Contract no. 2020-T1/TIC-19882 was granted by the Comunidad de Madrid in Spain, and the science research grants were from the China Manned Space Project. He also thanks the Ministerio de Ciencia e Innovación (Spain) for financial support under Project grant PID2021-122603NB-C21 and HORIZON EUROPE Marie Sklodowska-Curie Actions for supporting the LACEGAL-III project with grant number 101086388.This work has been made possible by the The Three Hundred collaboration [^1] https://www.the300-project.org. The HD simulations (7K and 15K runs) were performed on the MareNostrum Finisterrae3, and Cibeles Supercomputers through The Red Española de Supercomputación grants (AECT-2022-3-0027, AECT-2023-1-0013, AECT-2023-2-0004, AECT-2023-3-0023, AECT-2024-1-0026), on the DIaL3 – DiRAC Data Intensive service at the University of Leicester through the RAC15 grant: Seedcorn/ACTP317, and on the Niagara supercomputer at the SciNet HPC Consortium. DIaL3 is managed by the University of Leicester Research Computing Service on behalf of the STFC DiRAC HPC Facility (https://www.dirac.ac.uk). The DiRAC service at Leicester was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure. This work also used the DiRAC Complexity system, operated by the University of Leicester IT Services, which forms part of the STFC DiRAC HPC Facility (https://www.dirac.ac.uk). This equipment is funded by BIS National E-Infrastructure capital grant ST/K000373/1 and STFC DiRAC Operations grant ST/K0003259/1. DiRAC is part of the National e-Infrastructure. SciNet (Loken et al., 2010) is funded by Innovation, Science and Economic Development Canada; the Digital Research Alliance of Canada; the Ontario Research Fund: Research Excellence; and the University of Toronto.

### v2

@@ 6 Conclusions
    Sharper Stellar Truncation: The stellar orbiting profile exhibits a steeper truncation than the dark matter, consistent with a sharper splashback feature. This is in line with the findings of Deason et al. (2021) and O’Neil et al. (2021), and suggests that stellar material—in particular the ICL—may offer a more distinct observational signature of the splashback boundary.
@@
    Accretion Rate Dependence: Both components follow the same $r_{\rm t}$$\Gamma$ relation, enabling the use of stellar profiles as tracers of halo mass accretion history. This builds on the theoretical framework established by Adhikari et al. (2014), More et al. (2015), and Diemer (2017), and confirms that the splashback radius is sensitive to recent mass growth, with minimal dependence on halo mass or cosmology.
@@
    Observational Viability: Projected stellar density profiles allow recovery of the splashback radius $R_{\rm t}$ with modest scatter ($\sim 0.3R_{200m}$), demonstrating its observational viability. This complements observational detections of splashback features in galaxy density and weak lensing profiles (e.g. Chang et al. 2018a; Baxter et al. 2017; Gabriel-Silva and Sodré 2025) and supports the use of the ICL as a tracer of cluster boundaries, especially in deep imaging surveys.
@@
    Infalling and Inner Profile Correlations: The infalling stellar profiles correlate with those of the dark matter, and at high accretion rate the stellar profile reliably traces the dark matter near the scale radius $r_{s}$. This suggests that stellar material can be used to probe both the outer and inner dynamical structure of clusters, consistent with results from Pizzardo et al. (2024) and O’Neil et al. (2021).
@@
As upcoming surveys like Euclid, LSST, and Roman deliver deeper imaging and improved statistics, the stellar splashback radius may become a key tool in precision cluster cosmology. For this reason, there should be a focus on developing observational pipelines for ICL-based splashback detection, quantifying systematics due to projection and substructure, and integrating splashback measurements into multi-probe cosmological frameworks (e.g. Mpetha et al., 2025; Joshi et al., 2026).
@@ Acknowledgements
KW and CP acknowledge the support of the ARC Centre of Excellence for All Sky Astrophysics in 3 Dimensions (ASTRO 3D), through project number CE170100013. AK is supported by the Spanish Ministerio de Ciencia e Innovación, (MICINN) under research grant PID2021-122603NB-C21 as well as project PID2024-156100NB-C21 financed by MICIU /AEI/10.13039/501100011033/FEDER, UE. AK further thanks Slowdive again, this time for ’everything is alive’. WC is supported by the STFC AGP Grant ST/V000594/1, the Atracción de Talento Contract no. 2020-T1/TIC-19882 was granted by the Comunidad de Madrid in Spain, and the science research grants were from the China Manned Space Project. He also thanks the Ministerio de Ciencia e Innovación (Spain) for financial support under Project grant PID2021-122603NB-C21 and HORIZON EUROPE Marie Sklodowska-Curie Actions for supporting the LACEGAL-III project with grant number 101086388.This work has been made possible by the The Three Hundred collaboration [^1] https://www.the300-project.org. The HD simulations (7K and 15K runs) were performed on the MareNostrum Finisterrae3, and Cibeles Supercomputers through The Red Española de Supercomputación grants (AECT-2022-3-0027, AECT-2023-1-0013, AECT-2023-2-0004, AECT-2023-3-0023, AECT-2024-1-0026), on the DIaL3 – DiRAC Data Intensive service at the University of Leicester through the RAC15 grant: Seedcorn/ACTP317, and on the Niagara supercomputer at the SciNet HPC Consortium. DIaL3 is managed by the University of Leicester Research Computing Service on behalf of the STFC DiRAC HPC Facility (https://www.dirac.ac.uk). The DiRAC service at Leicester was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure. This work also used the DiRAC Complexity system, operated by the University of Leicester IT Services, which forms part of the STFC DiRAC HPC Facility (https://www.dirac.ac.uk). This equipment is funded by BIS National E-Infrastructure capital grant ST/K000373/1 and STFC DiRAC Operations grant ST/K0003259/1. DiRAC is part of the National e-Infrastructure. SciNet (Loken et al., 2010) is funded by Innovation, Science and Economic Development Canada; the Digital Research Alliance of Canada; the Ontario Research Fund: Research Excellence; and the University of Toronto.
@@
Appendix A Effect of the angular median on stellar profiles
Figure 13: Comparison between $r_{\rm s}$ and $\alpha$ obtained from fitting the dark matter and stellar orbiting profiles. The points are colored by accretion rate (Equation 4). The square points are the result of fitting the $\Gamma$-binned median profiles.
Figure 14: The same as Figure 12 but using angular-median density profiles.
As described in §3.2, the angular median method we use to calculate density profiles is much more effective at suppressing local fluctuations in density due to anisotropic substructure than the spherical average. This is because the median is much less sensitive to large outliers than the mean. Naturally, this affects the shape of the density profile and hence the parameters obtained from fitting.
Figure 13 compares the values of $r_{\rm s}$ and $\alpha$ determined from fitting the stellar and dark matter orbiting profiles. Unlike the parameters of the truncation term, $r_{\rm s}$ and $\alpha$ do not agree well between the stars and dark matter, despite what might be expected given the results presented in §4.4. However, this disagreement may be partly understood by looking at the equivalent of Figure 12 for the angular median profiles, shown in Figure 14. It is immediately clear that the density profile of stars in satellites is heavily suppressed—particularly around the dark matter $r_{\rm s}$. This is the expected effect of the angular median. However, the consequence is that the peak of the satellite profile is shifted and concealed by the BCG at all accretion rates, making the $r_{\rm s}$ measurement unreliable and disconnected from that of the dark matter. Visual inspection shows that this is reflected in the individual profiles, though there are exceptions.
Appendix B Satellite number density profiles
Figure 15: The mean stellar density slope profile of all GIZMO_7k clusters for several different definitions, where the density profiles are calculated using a spherical average (top) and the angular median described in §3.2 (bottom). The black curve is the total density of stars, the solid orange and blue curves are the density of particles contained inside satellites and outside satellites (i.e. BCG+ICL), respectively. The dashed orange curve is the density slope of satellites, calculated by treating the satellites as point particles with mass equal to their stellar mass (as defined by ahf). Finally, the orange dotted curve is the slope of the number density profile of satellites.
Given the challenges involved in observing the stellar density profile, it is worthwhile to assess how well we expect our results apply to galaxy number density profiles, which are more readily observable.
The top panel of Figure 15 compares the slope of the mean total mass density of all the GIZMO_7k clusters (black) to that of the number density of satellites $n_{\rm sat}$ (dotted orange), where the individual profiles have been calculated using a spherical average. We also show the slope of the density profile of particles in satellites (solid orange), of the smooth BCG+ICL (i.e. all other particles; blue), and of the density $\rho_{\rm sat}$ calculated by treating the satellites as point particles with mass equal to their stellar mass (as defined by ahf; dashed orange). All three profiles that are calculated using the satellites, regardless of the method, agree well with the total profile beyond $\sim 0.3R_{200{\rm m}}$. This is not unexpected given that the stellar mass outside the central region of clusters is dominated by stars bound in satellites. Meanwhile, the ICL profile is significantly different from the others, being steeper and having a slightly larger $r_{\rm steep}$. These results suggest that spherically-averaged satellite number density profiles accurately trace the total density profile around the splashback region.
The bottom panel of Figure 15 is the same as the top panel, except that the profiles being averaged are calculated using the angular median method described in §3.2. The angular median shifts $r_{\rm steep}$ slightly outwards but preserves the agreement between the profiles. It also makes the steepening feature of the profiles sharper, though this effect is not as significant for the number density. This is because the number density profiles effectively assign all satellites the same unit mass, which reduces the anistropy of the mass distribution in each radial shell. The angular median therefore has a significantly smaller effect on the number density than on the more anisotropic total mass density. However, this can (in principle) be mitigated by weighting the satellites by their stellar mass, as shown by the agreement between $\rho_{\rm sat}$ and total density slope around the splashback region. The sharpening of the splashback feature by the angular median may make it easier to identify in observations—even for clusters that don’t show a visible feature in the satellite number density.